A train begins at the origin and moves at constant speed v1=13.1 m/s toward another train which is initially at rest at a distance d=264 m from the first train. If the second train has an acceleration of magnitude a2=1.22 m/s2 toward the first train, how far does the first train move before the trains collide? (Hint : use quadratic equation and choose positive time)
Answer #1

47=1341 a=122 2 Si S2 26um 264 S + S₂ (uit) + [unt + £at²] = 264 13.1t lot 2x1-22+]=264 0.61t² + 13.1t - 264 = 0 + Ialo It =. Please like